# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ Appendix D Double-loop Glauber Dynamics on Weighted Graphs
$A_{i,j}=\begin{cases}w_{(i,j)}&\text{ if }(u_{i},u_{j})\in E;\\ 0&\text{ otherwise.}\end{cases}$
@@
$\operatorname{Haf}(G):=\operatorname{Haf}(A)=\frac{1}{2^{n}n!}\sum_{\sigma\in\mathcal{S}_{2n}}\prod_{i=1}^{n}A_{\sigma(2i-1),\sigma(2i)}.$
@@
$\operatorname{Haf}(G)=\sum_{M\text{ is a perfect matching of }G}\prod_{e\in M}w_{e}.$ (17)
@@
$\Pr[S]\propto\lambda^{|S|}\operatorname{Haf}^{2}(S).$
@@
2 Initialize $X_{0}$ as an arbitrary matching in $G$
3Initialize $t\leftarrow 0$, set the number of iterations $T$
@@ 2 Initialize $X_{0}$ as an arbitrary matching in $G$
5 Choose a uniformly random edge $e$ from $E$
@@
10Sample a perfect matching $E_{i}$ in the subgraph induced by $X_{t}$ with probability $\Pr[E_{i}]\propto w(E_{i})$. If $e\not\in E_{i}$, set $X_{t+1}=X_{t}$. If $e\in E_{i}$, set $X_{t+1}=X_{t}\backslash\{e\}$ with probability $\frac{1}{1+\lambda^{2}}\cdot\frac{1}{w_{e}^{2}}$ and otherwise set $X_{t+1}=X_{t}$;
@@ 14 Output the vertex set in matching $X_{T}$;
$\Pr[S]\propto\lambda^{|S|}\operatorname{Haf}^{2}(S).$
@@
$\Pr[X\to X\cup\{e\}]=\frac{\lambda^{2}}{1+\lambda^{2}}\cdot\frac{1}{|E|}.$
@@
$\Pr[e\in E_{i}]=\frac{\sum_{E^{\prime}\text{ is a perfect matching of }G_{X}}w(E^{\prime})\cdot w_{e}}{\sum_{E^{\prime}\text{ is a perfect matching of }G_{X\cup\{e\}}}w(E^{\prime})}=w_{e}\cdot\frac{\operatorname{Haf}(X)}{\operatorname{Haf}(X\cup\{e\})}.$
@@
$\Pr[X\cup\{e\}\to X]=\frac{1}{|E|}\cdot\frac{1}{w^{2}_{e}}\frac{1}{1+\lambda^{2}}\cdot w_{e}\frac{\operatorname{Haf}(X)}{\operatorname{Haf}(X\cup\{e\})},$ (18)
@@
$\pi(X)\propto\lambda^{2|E|}\cdot\operatorname{Haf}(G_{X})\cdot w(X),$
@@
$\pi(S)\propto\lambda^{|S|}\cdot\operatorname{Haf}(S)\cdot\sum_{\text{$X$ is a perfect matching of $G_{S}$}}w(X)=\lambda^{|S|}\cdot\operatorname{Haf}^{2}(S).$
Finally, we give a Markov chain to satisfy the desired sampling of perfect matchings in Line 5 in Algorithm 5. The method is based on the Markov chain for uniform sampling of perfect matchings by [24], which is discussed in Appendix B. The modified Markov chain is given as follows:
@@
  • If $M\in M_{n}(G)$ and $e\in M$, then move to state $M^{\prime}=M\backslash\{e\}$ with probability $\frac{1}{w_{e}}$;
@@
  • If $M\in M_{n-1}(G)$, $u$ is matched to $z$ in $M$ and $v$ is unmatched in $M$, then move to state $M^{\prime}=M\cup\{e\}\backslash\{(z,u)\}$ with probability $\min\{1,\frac{w_{e}}{w_{(z,u)}}\}$; symmetrically, if $v$ is matched to $z$ and $u$ is unmatched, then move to state $M^{\prime}=M\cup\{e\}\backslash\{(z,v)\}$ with probability $\min\{1,\frac{w_{e}}{w_{(z,v)}}\}$;
@@
For two near-perfect matchings $M$ and $M^{\prime}$ such that $M^{\prime}=M\cup\{(u,v)\}\backslash\{(z,u)\}$, the transition probability is given by
$\Pr[M\to M^{\prime}]=\frac{1}{|E|}\cdot\min\left\{1,\frac{w_{(u,v)}}{w_{(z,u)}}\right\};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}\cdot\min\left\{1,\frac{w_{(z,u)}}{w_{(u,v)}}\right\}.$
For a perfect matching $M$ and a near-perfect matching $M^{\prime}$ such that $M^{\prime}=M\backslash\{(u,v)\}$, the transition probability is given by
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}\cdot\frac{1}{w_{(u,v)}}.$
@@
Appendix E Algorithms with Pseudocodes
This section introduces the classical algorithms and their variants enhanced by Glauber dynamics we used for numerical experiments in Section V. During the iterations of the Glauber dynamics, we create a list. As long as the size of current edge set equals $k/2$, we append the edge set to the list. Each time we generate a vertex set of size $k$ according to the Glauber dynamics, we choose the last element in the list as the selected matching.
@@ Appendix E Algorithms with Pseudocodes
1 Initialize $i=0$, $Best=0$;
2Initialize $\bm{A}_{B}=None$;
3while $i<N$ do
@@ 4 Choose $\bm{A}_{i}$, a $k$-dimensional submatrix of $\bm{A}$ uniformly at random;
5Calculate $f(\bm{A}_{i})$;
6if $f(\bm{A}_{i})>Best$ then
@@ 8    $\bm{A}_{B}=\bm{A}_{i}$;
1 Initialize $i=0$, $Best=0$;
2Initialize $\bm{A}_{B}=None$;
3while $i<N$ do
@@ 1 Initialize $i=0$, $Best=0$;
5Calculate $f(\bm{A}_{i})$;
6if $f(\bm{A}_{i})>Best$ then
@@ 8    $\bm{A}_{B}=\bm{A}_{i}$;
Instead of applying i.i.d. samples in random search, we adopt a neighbor update strategy in simulated annealing (SA). In each iteration, we randomly generate an integer $m\in[0,k-1]$ and randomly choose $m$ of the former vertices to keep unchanged. Then we randomly sample $k-m$ vertices from the remaining $n-m$ vertices to form a new vertex set of size $k$, and update the set with probability assigned by a Metropolis filter. With the annealing parameter $10<\gamma<$ 1, the temperature approaches exponentially small as iteration increases and the update probability when $f(\bm{A}_{R})<(\bm{A}_{S})$ approaches 00. For complex search space with large numbers of local optima metaheuristic SA is capable of approximating the global optimum, due to its probabilistic acceptance of worse solutions which enables exploration beyond local minima [31]. The pseudocode is presented in Algorithm 8.
@@
2 Generate a uniformly random binary vector $S$ of length $n$ with $k$ entries being one;
3Get the $k$-dimensional submatrix $\bm{A}_{S}$ according to $S$;
4Initialize $\bm{A}_{B}=\bm{A}_{S}$ and $Best=f(\bm{A}_{S})$;
5Initialize $i=0$;
6while $i<N$ do
@@ 7 Randomly generate an integer $m\in[0,k-1]$;
8Randomly choose $m$ nonzero entries of $S$ and keep them unchanged, then resample $k-m$ nonzero entries from the remaining $n-m$ vertices to get $R$;
9Get $\bm{A}_{R}$ according to $R$;
10Calculate $f(\bm{A}_{R})$;
11if $f(\bm{A}_{R})>Best$ then
@@
14Set $S=R$ with probability $\min\{1,\exp\left[\frac{f\left(\bm{A}_{R}\right)-f\left(\bm{A}_{S}\right)}{t}\right]\}$
@@ 15   $t~{}*=\gamma$
2 Generate a binary vector $S$ of length $n$ with $k$ entries being one according to Algorithm 1 or Algorithm 2 or Algorithm 3;
3Get a $k$-dimensional submatrix $\bm{A}_{S}$ according to $S$;
4Initialize $\bm{A}_{B}=\bm{A}_{S}$ and $Best=f(\bm{A}_{S})$;
5Initialize $i=0$;
6while $i<N$ do
@@ 7 Randomly generate an integer $m\in[0,k-1]$;
8Generate $s$ by randomly choosing $m$ nonzero entries of $S$ and set the rest entries as zeroes;
9Get $\bm{A}_{{R}}$ according to $R$;
10Generate $r$ with $k$ entries are nonzero according to Algorithm 1 or Algorithm 2 or Algorithm 3 and set $m$ entries as zeroes randomly under the condition that there are no overlap between nonzero elements of $s$ and $r$;
[…]
12Calculate $f(\bm{A}_{R})$;
13if $f(\bm{A}_{R})>Best$ then
@@
16Set $S=R$ with probability $\min\{1,\exp\left[\frac{f\left(\bm{A}_{R}\right)-f\left(\bm{A}_{S}\right)}{t}\right]\}$

### v2

@@ E Technical Barrier for Non-dense Graphs
$A_{i,j}=\begin{cases}w_{(i,j)}&\text{ if }(u_{i},u_{j})\in E;\\ 0&\text{ otherwise.}\end{cases}$ (S94)
@@
$\operatorname{Haf}(G):=\operatorname{Haf}(A)=\frac{1}{2^{n}n!}\sum_{\sigma\in\mathcal{S}_{2n}}\prod_{i=1}^{n}A_{\sigma(2i-1),\sigma(2i)}.$ (S95)
@@
$\operatorname{Haf}(G)=\sum_{M\text{ is a perfect matching of }G}\prod_{e\in M}w_{e}.$ (S96)
@@
$\Pr[S]\propto\lambda^{|S|}\operatorname{Haf}^{2}(S).$ (S97)
@@
2 Initialize $X_{0}$ as an arbitrary matching in $G$;
3Initialize $t\leftarrow 0$, set the number of iterations $T$;
@@ 2 Initialize $X_{0}$ as an arbitrary matching in $G$;
5 Choose a uniformly random edge $e$ from $E$;
@@
[…]
@@ 14 Output the vertex set in matching $X_{T}$;
$\Pr[S]\propto\lambda^{|S|}\operatorname{Haf}^{2}(S).$ (S98)
@@
$\Pr[X\to X\cup\{e\}]=\frac{\lambda^{2}}{1+\lambda^{2}}\cdot\frac{1}{|E|}.$ (S99)
@@
[…]
@@
$\Pr[X\cup\{e\}\to X]=\frac{1}{|E|}\cdot\frac{1}{w^{2}_{e}}\frac{1}{1+\lambda^{2}}\cdot w_{e}\frac{\operatorname{Haf}(X)}{\operatorname{Haf}(X\cup\{e\})},$ (S101)
@@
$\pi(X)\propto\lambda^{2|E|}\cdot\operatorname{Haf}(G_{X})\cdot w(X),$ (S102)
@@
$\pi(S)\propto\lambda^{|S|}\cdot\operatorname{Haf}(S)\cdot\sum_{\text{$X$ is a perfect matching of $G_{S}$}}w(X)=\lambda^{|S|}\cdot\operatorname{Haf}^{2}(S).$ (S103)
Finally, we give a Markov chain to satisfy the desired sampling of perfect matchings in Line 5 in Algorithm 5. The method is based on the Markov chain for uniform sampling of perfect matchings by [31], which is discussed in Supplementary Note 2. The modified Markov chain is given as follows:
@@
  • If $M\in M_{n}(G)$ and $e\in M$, then move to state $M\{$$^{\prime}=M\backslash\{e\}$ with probability $\frac{1}{w_{e}}$;
@@
[…]
@@
[…]
$\Pr[M\to M^{\prime}]=\frac{1}{|E|}\cdot\min\left\{1,\frac{w_{(u,v)}}{w_{(z,u)}}\right\};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}\cdot\min\left\{1,\frac{w_{(z,u)}}{w_{(u,v)}}\right\}.$ (S105)
[…]
$\Pr[M\to M^{\prime}]=\frac{1}{|E|};\quad\Pr[M^{\prime}\to M]=\frac{1}{|E|}\cdot\frac{1}{w_{(u,v)}}.$ (S106)
@@
Supplementary Note 5 Algorithms with Pseudocodes
This section introduces the classical algorithms and their variants enhanced by Glauber dynamics we used for numerical experiments in Section II.4. In the actual code execution, we use the method in Supplementary Note A to uniformly sample perfect matching. During the iterations of the Glauber dynamics, we create a post-selection list. As long as the size of current edge set equals $k/2$, we append the edge set to the list. Each time we generate a vertex set of size $k$ according to the Glauber dynamics, we choose the last element in the list as the selected matching.
@@
1
21em Initialize $i=0$, $Best=0$;
3Initialize $\bm{A}_{B}=None$;
4while $i<N$ do
@@ 5 Choose $\bm{A}_{i}$, a $k$-dimensional submatrix of $\bm{A}$ uniformly at random;
6Calculate $f(\bm{A}_{i})$;
7if $f(\bm{A}_{i})>Best$ then
@@ 9    $\bm{A}_{B}=\bm{A}_{i}$;
1
21em Initialize $i=0$, $Best=0$;
3Initialize $\bm{A}_{B}=None$;
4while $i<N$ do
@@
6Calculate $f(\bm{A}_{i})$;
7if $f(\bm{A}_{i})>Best$ then
@@
Instead of applying i.i.d. samples in random search, we adopt a neighbor update strategy in simulated annealing (SA). In each iteration, we randomly generate an integer $m\in[0,k-1]$ and randomly choose $m$ of the former vertices to keep unchanged. Then we randomly sample kmk-m vertices from the remaining nmn-m vertices to form a new vertex set of size $k$, and update the set with probability assigned by a Metropolis filter. With the annealing parameter $0<\gamma<1$, the temperature approaches exponentially small as iteration increases and the update probability when $f(\bm{A}_{R})<(\bm{A}_{S})$ approaches 0. For complex search space with large numbers of local optima metaheuristic SA is capable of approximating the global optimum, due to its probabilistic acceptance of worse solutions which enables exploration beyond local minima [42]. The pseudocode is presented in Algorithm 8.
@@
2
31em Generate a uniformly random binary vector $S$ of length $n$ with $k$ entries being one;
4Get the $k$-dimensional submatrix $\bm{A}_{S}$ according to $S$;
5Initialize $\bm{A}_{B}=\bm{A}_{S}$ and $Best=f(\bm{A}_{S})$;
6Initialize $i=0$;
7while $i<N$ do
@@ 8 Randomly generate an integer $m\in[0,k-1]$;
9Randomly choose $m$ nonzero entries of $S$ and keep them unchanged, then resample kmk-m nonzero entries from the remaining nmn-m vertices to get $R$;
10Get $\bm{A}_{R}$ according to $R$;
11Calculate $f(\bm{A}_{R})$;
12if $f(\bm{A}_{R})>Best$ then
@@
15Set $S=R$ with probability $\min\{1,\exp\left[\frac{f\left(\bm{A}_{R}\right)-f\left(\bm{A}_{S}\right)}{t}\right]\}$
@@ 16   $t~*=\gamma$
2
31em Generate a binary vector $S$ of length $n$ with $k$ entries being one according to Algorithm 1 or Algorithm 2 or Algorithm 3;
4Get a $k$-dimensional submatrix $\bm{A}_{S}$ according to $S$;
5Initialize $\bm{A}_{B}=\bm{A}_{S}$ and $Best=f(\bm{A}_{S})$;
6Initialize $i=0$;
7while $i<N$ do
@@ 8 Randomly generate an integer $m\in[0,k-1]$;
9Generate $s$ by randomly choosing $m$ nonzero entries of $S$ and set the rest entries as zeroes;
10Get $\bm{A}_{{R}}$ according to $R$;
11Generate $r$ with $k$ entries are nonzero according to Algorithm 1 or Algorithm 2 or Algorithm 3 and set $m$ entries as zeroes randomly under the condition that there are no overlap between nonzero elements of $s$ and $r$;
12R=s|rR=s|r;
13Calculate $f(\bm{A}_{R})$;
14if $f(\bm{A}_{R})>Best$ then
@@
17Set $S=R$ with probability $\min\{1,\exp\left[\frac{f\left(\bm{A}_{R}\right)-f\left(\bm{A}_{S}\right)}{t}\right]\}$